A square pizza is sold for $15.00 and is 10 inches on each side. Using the price per square inch of the square piza, determine a fair price for a circular pizza that has a diameter of 10 inches. Show how you found the price per square inch of the square pizza, the area of the circular pizza, and eplain how
you used both pieces of information to determine a fair price for the circular pizza.

Answer by CharlesG2(834)

not Geometry: Volume, Metric volume
moved it to Geometry: Area and Surface Area
"A square pizza is sold for $15.00 and is 10 inches on each side. Using the price per square inch of the square piza, determine a fair price for a circular pizza that has a diameter of 10 inches. Show how you found the price per square inch of the square pizza, the area of the circular pizza, and eplain how
you used both pieces of information to determine a fair price for the circular pizza."
Area of square pizza = 10 inches * 10 inches = 100 square inches (sq.in.)
Price per square inch of square pizza = $15.00 / 100 sq.in. = $0.15 per sq.in.
Area circular pizza = pi * r^2, where r is the radius
r = diameter/2 = 10/2 inches = 5 inches
Area circular pizza = pi * 5^2 = 25pi sq.in.
Price circular pizza = 25pi * 1 sq.in. * $0.15 * 1/sq.in.
Price circular pizza = 25pi * $0.15 = $3.75pi
Price circular pizza = approx. $11.78